# Stack More Layers Differently: High-Rank Training Through Low-Rank Updates

### v1

@@ Abstract
Despite the dominance and effectiveness of scaling, resulting in large networks with hundreds of billions of parameters, the necessity to train overparametrized models remains poorly understood, and alternative approaches do not necessarily make it cheaper to train high-performance models. In this paper, we explore low-rank training techniques as an alternative approach to training large neural networks. We introduce a novel method called ReLoRA, which utilizes low-rank updates to train high-rank networks. We apply ReLoRA to pre-training transformer language models with up to 350M parameters, and demonstrate comparable performance to regular neural network training. Furthermore, we observe that the efficiency of ReLoRA increases with model size, making it a promising approach for training multi-billion-parameter networks efficiently. Our findings shed light on the potential of low-rank training techniques and their implications for scaling laws. Code is available on GitHub.111github.com/guitaricet/peft_pretraining
@@ 1 Introduction
Over the past decade, the machine learning field has been dominated by the trend of training increasingly overparametrized networks or adopting the "stack more layers" approach Krizhevsky et al. (2012); He et al. (2016); Kaplan et al. (2020). The definition of a large network has evolved from models with 100 million Simonyan and Zisserman (2015); Radford et al. (2018) to hundreds of billions Brown et al. (2020); Chowdhery et al. (2022) of parameters, which has made computational costs associated with training of such networks prohibitive to most of the research groups. Despite this, the necessity to train models which can have orders of magnitude more parameters than the training examples Brown et al. (2020); Chowdhery et al. (2022); Fedus et al. (2022), is poorly understood theoretically Jacot et al. (2018); Allen-Zhu et al. (2019); Zhang et al. (2021).
Alternative approaches to scaling, such as more compute-efficient scaling optima Hoffmann et al. (2022), retrieval-augmented models Khandelwal et al. (2020); Borgeaud et al. (2022), and the simple approach of training smaller models for longer Touvron et al. (2023), have offered new interesting trade-offs. However, they do not bring us closer to understanding why we need overparametrized models and rarely democratize the training of these models. For example, training RETRO Borgeaud et al. (2022) requires a complex training setup and infrastructure capable of quickly searching over trillions of tokens, while training LLaMA-6B Touvron et al. (2023) still requires hundreds of GPUs.
@@
On the one hand, pre-training is exactly the step that allows for small modifications to the network to adapt it to new tasks. Aghajanyan et al. (2021) demonstrated that the rank of the changes required to learn a task decreases the more you pre-train the network. On the other hand, multiple studies have demonstrated the simplicity of features extracted and utilized by language and vision models, along with their low intrinsic dimensionality Kovaleva et al. (2019); Frankle and Carbin (2019); Raganato et al. (2020); Singh et al. (2021). For instance, attention patterns in transformers Vaswani et al. (2017) often exhibit a small rank, which has been successfully leveraged to develop more efficient variants of attention Wang et al. (2020); Choromanski et al. (2021). Moreover, overparametrization is also not necessary for training. The Lottery Ticket Hypothesis Frankle and Carbin (2019) empirically demonstrates that during initialization (or early in training Frankle et al. (2019)), there exist sub-networks – winning tickets – that when trained in isolation reach the performance of the full network.
@@
In this study, we focus on low-rank training techniques and introduce ReLoRA that uses low-rank updates to train a high-rank network. We empirically demonstrate that ReLoRA performs a high-rank update and achieves performance similar to regular neural network training. The components of ReLoRA include initial full-rank training of the neural network (similar to Frankle et al. (2019)), LoRA training, restarts, a jagged learning rate schedule, and partial optimizer resets. We evaluate ReLoRA on transformer language models up to 350M parameters. We chose to focus on autoregressive language modeling, as this approach has demonstrated its universality in most of the applications of neural networks Raffel et al. (2020); Yu et al. (2022); Alayrac et al. (2022); Lu et al. (2023); Cho et al. (2021). Finally, we observe that the efficiency of ReLoRA increases with model size, making it a viable option for efficient training of multi-billion-parameter networks.
Each experiment in this study has used no more than 8 GPU days of compute.
@@
Figure 1: ReLoRA learns a high-rank network through a sequence of low-rank updates. It outperforms networks with the same trainable parameter count and achieves similar performance to training a full network at 100M+ scale. The efficiency of ReLoRA increases with the model size, making it a viable candidate for multi-billion-parameter training.
2 Related work
Scaling versus Efficiency
The relationship between overparametrization and neural network trainability and generalization has been extensively studied Zhang et al. (2017); Belkin et al. (2018); Frankle and Carbin (2019); Nakkiran et al. (2019); Singh et al. (2021), yet it remains a mystery Zhang et al. (2021). Moreover, scaling laws Kaplan et al. (2020); Ghorbani et al. (2021); Hoffmann et al. (2022); Klug and Heckel (2023); Aghajanyan et al. (2023) demonstrate a simple and strong power-law dependence between network size and its performance across a variety of modalities. This finding not only supports overparametrization but also encourages the training of extraordinarily resource-intensive neural networks Brown et al. (2020); Chowdhery et al. (2022); Fedus et al. (2022). Nonetheless, the Lottery Ticket Hypothesis Frankle and Carbin (2019); Frankle et al. (2019) suggests that overparametrization could, in principle, be minimized. Specifically, it shows that early in training, subnetworks exist that can be trained to achieve the performance of the full network (winning tickets).
Parameter-efficient fine-tuning
Aghajanyan et al. (2021) found that pre-training reduces the amount of change to the network, or its intrinsic dimensionality, to learn a new task through fine-tuning. I.e., larger networks or networks pre-trained on more data require smaller modifications in terms of the rank of the range to learn a new task. This explains the success of parameter-efficient fine-tuning methods Lialin et al. (2023) and has also motivated the development of low-rank fine-tuning methods such as LoRA Hu et al. (2022) and Compacter mahabadi et al. (2021).
Low-rank neural network training
Training low-rank representations has been explored in the context of CNN compression, regularization, and efficient training Idelbayev and Carreira-Perpinan (2020); Jaderberg et al. (2014); Sui et al. (2023); Schotthöfer et al. (2022); Lin et al. (2020); Yuan et al. (2021). However, most of these methods are either specific to CNNs, do not scale well, or have not been evaluated on large transformers Vaswani et al. (2017) with hundreds of millions of parameters, which can benefit greatly from efficient training. While transformers have been shown to have a low-rank internal dimensionality and representations Aghajanyan et al. (2021); Wang et al. (2020), the study by Bhojanapalli et al. (2020) demonstrated that the low rank of key and query projections in multi-head attention bottlenecks the performance of transformers. Our own experiments (Section 3) also demonstrate that low-rank transformers perform significantly worse compared to the full-rank baseline and ReLoRA.
@@ 3 Method
Let’s start by revisiting linear algebra-101. In particular, we are interested in the rank of the sum of two matrices:
$\text{rank}(A+B)\leq\text{rank}(A)+\text{rank}(B).$ (1)
This bound on the rank of the sum is tight: for a matrix $\mathbf{A},\text{rank}(\mathbf{A})<dim(\mathbf{A})$, there exists $\mathbf{B}$, $\text{rank}(\mathbf{B})<dim(\mathbf{B})$ such that sum of the matrices has a higher rank than either $\mathbf{A}$ or $\mathbf{B}$. We want to exploit this property to make a flexible parameter-efficient training method. We start with LoRA Hu et al. (2022) which is a parameter-efficient fine-tuning method based on the idea of low-rank updates. LoRA can be applied to any linear operation parametrized through $W\in\mathbb{R}^{m\times n}$. Specifically, LoRA decomposes the weight update $\delta W$ into a low-rank product $W_{A}W_{B}$ as shown in Equation 2, where $s\in\mathbb{R}$ is a fixed scaling factor usually equal to $\frac{1}{r}$.
@@
In practice, LoRA is usually implemented by adding new trainable parameters $W_{A}$ and $W_{B}$, which could be merged back into the original parameters after training. Thus, even though Equation 1 allows the total update over training time $\sum_{t}\delta W_{t}$ to have a higher rank than any of the individual matrices, LoRA implementations are restricted by the rank $r=\max_{W_{A},W_{B}}\text{rank}(W_{A}W_{B})$.
If we could restart LoRA, meaning we merge $W_{A}$ and $W_{B}$ during training and reset the values of these matrices, we could increase the total rank of the update. Doing this multiple times brings the total neural network update to
Figure 2: Jagged cosine scheduler used in ReLoRA. On every ReLoRA reset, we set the learning rate to zero and perform a quick (50-100 steps) learning rate warmup back to the cosine schedule.
@@
where the sums are independent enough, meaning that $\text{rank}(W_{A}^{i}W_{B}^{i})+\text{rank}(W_{A}^{j}W_{B}^{j})\geq r$.
However, implementing restarts is not trivial in practice and requires certain modifications to the optimization procedure. Naïve implementation causes the model to diverge right after the restart. Unlike plain stochastic gradient descent, which solely relies on the value of the gradient at the current optimization timestep, Adam Kingma and Ba (2015) update is guided mainly by the first and second moments of the gradient accumulated over the previous steps. In practice, gradient moment smoothing parameters $\beta_{1}$ and $\beta_{2}$ are usually very high $0.9-0.999$. Let’s assume that at the reinitialization boundary $W_{A}^{1}$ and the corresponding gradient moments $m_{A}$ and $v_{A}$, are full-rank ($r$). Then, after the merge-and-reinit, continuing to use old gradient moments for $W_{A}^{2}$ will guide it in the same direction as $W_{A}^{1}$ and optimize the same subspace.
To resolve this issue, we propose ReLoRA. ReLoRA performs a partial reset of the optimizer state during merge-and-reinit and sets the learning rate to 0 with a subsequent warmup. Specifically, we set $99\%$ of low-magnitude optimizer state values to zero and use a jagged-cosine learning rate schedule (Figure 2). Our ablation studies (Section 3) show that both of these modifications are required to improve the performance over vanilla LoRA.
To reiterate, ReLoRA is a low-rank training method inspired by LoRA that uses restarts to increase the effective rank of the update, uses partial optimizer reset, and a jagged scheduler to stabilize training and warm starts. All of this allows ReLoRA to achieve performance comparable to full-rank training, especially in large transformer networks, by only training a small set of parameters at a time. ReLoRA is described in Algorithm 1.
@@
Unlike other low-rank training techniques Schotthöfer et al. (2022); Sui et al. (2023), ReLoRA follows the LoRA approach by maintaining the frozen weights of the original network and adding new trainable parameters. At first glance, this may appear computationally inefficient; however, the differentiation between frozen and trainable parameters plays a crucial role in parameter-efficient fine-tuning Lialin et al. (2023).
These methods achieve significant improvements in training time and memory efficiency by reducing the size of the gradients and the optimizer states. Notably, Adam states consume twice as much memory as the model weights. Moreover, it is common practice to maintain gradient accumulation buffers in 32-bit precision for large networks, thereby adding significant overhead to the memory consumption of gradients.
By substantially reducing the number of trainable parameters, ReLoRA enables the utilization of larger batch sizes, maximizing hardware efficiency. Additionally, it reduces the bandwidth requirements in distributed setups, which are often the limiting factor in large-scale training.
Furthermore, since the frozen parameters are not being updated between restarts, they can be kept in a low-precision quantized format, further reducing their memory and computational impact. This additional optimization contributes to overall improved efficiency in terms of memory utilization and computational resources of ReLoRA and increases at scale.
@@ 4 Experiments
To evaluate the effectiveness of ReLoRA, we apply it to train a transformer language model on the C4 dataset Raffel et al. (2020) using various model sizes: 60M, 130M, 250M, and 350M. Language modeling has been shown to be a fundamental task in machine learning Radford et al. (2019), it enables text and image classification Yu et al. (2022), translation Brown et al. (2020), programming Chen et al. (2021), in-context learning, step-by-step reasoning Wei et al. (2022b), and many other emergent abilities Wei et al. (2022a). Given its significance, we focus solely on language modeling for the purposes of this paper.
Algorithm 1 ReLoRA. $\theta$ is model parameters, $\hat{\theta}$ is model parameters with linear layers replaced with ReLoRA, $M$ and $V$ are Adam optimizer states, $\eta$ is learning rate scheduled according to a jagged scheduler, and finally, $q$ is the reinit frequency.
@@
Params Hidden Heads Layers Learning rate Batch (tokens) Seq. len. Tokens
@@
Table 1: Hyperparameters of the language models trained in this study.
@@
Our architecture is based on transformer Vaswani et al. (2017) and closely resembles LLaMA Touvron et al. (2023). Namely, we use pre-normalization, RMSNorm Zhang and Sennrich (2019), SwiGLU activations Shazeer (2020), $\frac{8}{3}h$ fully-connected hidden state size Touvron et al. (2023), and rotary embeddings Su et al. (2021). All hyperparameters are presented in Table 1.
We use bfloat16 for all floating point operations and Flash attention Dao et al. (2022) for effective attention computation. Compared to attention in LLaMA, which uses float32 for softmax computation, this increased training throughput by 50-100% without any training stability issues.
Most of our models were trained on 8 RTX 4090 for one day or less. Due to computational constraints, we train much smaller models than LLaMA, with the largest model having 350M parameters, the same as BERT Large Devlin et al. (2019). We select the number of pre-training tokens based on the Chinchilla scaling laws Hoffmann et al. (2022) for all models, except for the largest one, which we train for 6.8B tokens while 9.5B tokens are Chinchilla-optimal.
@@
In our low-rank training experiments, ReLoRA replaces all attention and fully-connected network parameters, while keeping the embeddings full-rank. The RMSNorm parametrization remains unchanged. Since ReLoRA-wrapped models have fewer trainable parameters than full-rank training, we include a Control baseline, which is a full-rank transformer with the same number of trainable parameters as ReLoRA.
We initialize ReLoRA from a checkpoint of full-rank training at 5,000 update steps and reset it every 5,000 steps thereafter, 3 times in total. After each reset, 99% of the optimizer state is pruned based on magnitude, and the loss is warmed up for the next 100 iterations. ReLoRA parameters are reinitialized following LoRA best practices, Kaiming initialization He et al. (2015) for $A$-matrix, and zeros for $B$-matrix. In case of not using the restarts, the $B$-matrix also uses Kaiming initialization to avoid gradient-symmetry issues.
@@ 5 Results
Our main results are resented in Table 2. ReLoRA significantly outperforms low-rank LoRA training demonstrating the effectiveness of our proposed modifications (ablated in Section 3). Furthermore, ReLoRA achieves similar performance to full-rank training, and the performance gap diminishes as network size increases.
Interestingly, the only model in which ReLoRA couldn’t surpass the Control baseline was our smallest model with 60M parameters. This observation suggests that ReLoRA is particularly effective in improving the training of large networks, which aligns with our goal of developing a method that improves large-network training.
@@
To determine whether ReLoRA performs a higher rank update than LoRA we plot the singular value spectrum of the difference between warm-start weights and the final weights for ReLoRA, LoRA, and full-rank training. Figure 3 illustrates significant qualitative differences between LoRA and ReLoRA for the singular values of $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{down}$.
While most of the singular values for LoRA are zero (Figure 4) with a noticeable number of exceptionally high values above 1.5, ReLoRA exhibits a higher distribution mass between 0.1 and 1.0, reminiscent of full-rank training. This observation emphasizes the significance of high-rank updates and demonstrates the qualitative efficacy of ReLoRA, which accomplishes a high-rank update by performing multiple low-rank updates.
60M 130M 250M 350M
Full training 33.81 23.65 22.39 20.40
Control 36.52 27.30 29.12 23.65
Low-rank pre-training with LoRA 47.44 34.17 36.60 57.11
Low-rank pre-training with ReLoRA 38.28 25.04 22.48 22.48
No. of training tokens (billions) 1.2 2.6 6.8 6.8
Table 2: Comparing perplexities between baseline methods and ReLoRA (lower is better). Control has the same number of trainable parameters as low-rank training. Low-rank training is bold if it outperforms the Control baseline. ReLoRA efficacy increases as the network size grows. NOTE: 350M run is WIP.
@@
Figure 3: Singular values spectra of the weight difference between ReLoRA and LoRA at 5,000 iterations (warm start) and 20,000 iterations. ReLoRA exhibits a closer resemblance to full-rank training singular values than LoRA, indicating its effectiveness in approximating full-rank behavior.
Restarts Jagged Schedule Optimizer Reset Warm Start High LR Perplexity ($\downarrow$)
$\times$ $\times$ $\times$ $\times$ $\times$ 34.17
$\checkmark$ $\times$ $\times$ $\times$ $\times$ 34.25
$\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ N/A
$\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 34.29
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 29.77
$\times$ $\times$ $\times$ $\checkmark$ $\times$ 25.46
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 25.04
Table 3: Ablation studies of ReLoRA. Restarts and warm starts are essential for good performance. Using restarts and optimizer reset without a jagged schedule causes the model to diverge.
@@
Figure 4: Number of singular values $<0.1$ in attention and FCN projection matrices.

### v4

# ReLoRA: High-Rank Training Through Low-Rank Updates

@@ Abstract
Despite the dominance and effectiveness of scaling, resulting in large networks with hundreds of billions of parameters, the necessity to train overparameterized models remains poorly understood, while training costs grow exponentially. In this paper, we explore parameter-efficient training techniques as an approach to training large neural networks. We introduce a novel method called ReLoRA, which utilizes low-rank updates to train high-rank networks. We apply ReLoRA to training transformer language models with up to 1.3B parameters and demonstrate comparable performance to regular neural network training. ReLoRA saves up to 5.5Gb of RAM per GPU and improves training speed by 9-40% depending on the model size and hardware setup. Our findings show the potential of parameter-efficient techniques for large-scale pre-training. Our code is available on GitHub111github.com/guitaricet/relora.
@@ 1 Introduction
Over the past decade, the machine learning field has been dominated by the trend of training increasingly overparameterized networks or adopting the "stack more layers" approach Krizhevsky et al. (2012); He et al. (2016); Kaplan et al. (2020). The definition of a large network has evolved from models with 100 million Simonyan and Zisserman (2015); Radford et al. (2018) to hundreds of billions Brown et al. (2020); Chowdhery et al. (2022) of parameters, which has made computational costs associated with training of such networks prohibitive to most of the research groups. Despite this, the necessity to train models which can have orders of magnitude more parameters than the training examples Brown et al. (2020); Chowdhery et al. (2022); Fedus et al. (2022), is poorly understood theoretically Jacot et al. (2018); Allen-Zhu et al. (2019); Zhang et al. (2021).
Alternative approaches to scaling, such as more compute-efficient scaling optima Hoffmann et al. (2022), retrieval-augmented models Khandelwal et al. (2020); Borgeaud et al. (2022), and the simple approach of training smaller models for longer Touvron et al. (2023), have offered new trade-offs. However, they do not bring us closer to understanding why we need overparameterized models and rarely democratize the training of these models. For example, training RETRO Borgeaud et al. (2022) requires a complex training setup and infrastructure capable of quickly searching over trillions of tokens, while training LLaMA-7B Touvron et al. (2023) still requires hundreds of GPUs.
@@
Our Contribution
@@
In this study, we introduce ReLoRA which uses individually low-rank updates that aggregate during the training process to train a high-rank network. We empirically demonstrate that ReLoRA performs a high-rank update and achieves performance similar to regular neural network training. The components of ReLoRA include initial full-rank training of the neural network (similar to Frankle et al. (2019)), LoRA training, restarts, a jagged learning rate schedule, and partial optimizer resets. We evaluate ReLoRA on transformer language models up to 1.3B parameters. Finally, we observe that the efficiency of ReLoRA increases with model size, making it a viable option for efficient training of multi-billion-parameter networks.
@@
Figure 1: Training loss for 250M models. ReLoRA learns a high-rank network through a sequence of low-rank updates. It outperforms networks with the same trainable parameter count and achieves similar performance to training a full network at 100M+ scale. The efficiency of ReLoRA increases with the model size, making it a viable candidate for multi-billion-parameter training.
@@ 2 Method
We are interested in the rank of the sum of two matrices: $\text{rank}(A+B)\leq\text{rank}(A)+\text{rank}(B)$. We know that for a matrix $\mathbf{A},\text{rank}(\mathbf{A})<dim(\mathbf{A})$, there exists a $\mathbf{B}$, $\text{rank}(\mathbf{B})<dim(\mathbf{B})$ such that sum of them has a higher rank than either $\mathbf{A}$ or $\mathbf{B}$.
We want to exploit this property to make a flexible parameter-efficient training method. We start with LoRA Hu et al. (2022) which is a parameter-efficient fine-tuning method based on the idea of low-rank updates. LoRA can be applied to any linear operation parametrized through $W\in\mathbb{R}^{m\times n}$. Specifically, LoRA decomposes the weight update $\delta W$ into a rank-$r$ product $W_{A}W_{B}$ as shown in Equation 1, where $s\in\mathbb{R}$ is a fixed scaling factor usually equal to $\frac{1}{r}$.
@@
In practice, LoRA is usually implemented by adding new trainable parameters $W_{A}$ and $W_{B}$, which could be merged back into the original parameters after training. Thus, these implementations are restricted by the rank $r=\max_{W_{A},W_{B}}\text{rank}(W_{A}W_{B})$.
If we could restart LoRA, meaning we merge $W_{A}$ and $W_{B}$ during training and reset the values of these matrices, we could increase the total rank of the update. Doing this multiple times brings the total neural network update to:
Figure 2: Jagged cosine scheduler used in ReLoRA. As a base for our scheduler we follow a standard cosine decay schedule as in Touvron et al. (2023). On every optimizer reset, we set the learning rate to zero and perform a quick (50-100 steps) learning rate warm-up back to the cosine schedule.
@@
However, implementing restarts is not trivial in practice and requires several modifications to the optimization procedure. Unlike plain stochastic gradient descent, Adam Kingma and Ba (2015) update is guided mainly by the first and second moments of the gradient accumulated over the previous steps. In practice, Adam’s $\beta_{1}$ and $\beta_{2}$ are usually very high $0.9-0.999$. This means that after the merge-and-reinit, continuing to use old gradient moments for $W_{A}^{2}$ will guide it in the same direction as $W_{A}^{1}$ and optimize the same subspace.
To resolve this issue, ReLoRA performs a partial reset of the optimizer state during merge-and-reinit via magnitude pruning. To avoid loss diverging after an optimizer reset it also sets the learning rate to 0 with a subsequent warm-up (Figure 2). Our ablation studies (Table 6) show that both of these modifications are required to improve the performance over LoRA. Finally, in our experiments we found that in the case of training from scratch (random initialization) a short full-rank training is needed to “warm start” ReLoRA. All of this allows ReLoRA to achieve performance comparable to full-rank training, especially in large transformer networks, by only training a small set of parameters at a time. ReLoRA is described in Algorithm 1.
@@
Unlike other low-rank training techniques Schotthöfer et al. (2022); Sui et al. (2023); Kamalakara et al. (2022), ReLoRA follows the LoRA approach by maintaining the frozen weights of the original network and adding new trainable parameters. At first glance, this may appear computationally inefficient; however, the differentiation between frozen and trainable parameters plays a crucial role in parameter-efficient fine-tuning Lialin et al. (2023).
By reducing the number of trainable parameters, ReLoRA significantly reduces the memory spent on the optimizer states and enables the utilization of larger batch sizes, maximizing hardware efficiency. Additionally, it reduces the bandwidth requirements in distributed setups, which are often the limiting factor in large-scale training. Furthermore, since the frozen parameters are not being updated between restarts, they can be kept in a low-precision quantized format Dettmers et al. (2023), further reducing their memory and computational impact.
Locally Low-Rank Training: Intuition
Multiple studies suggest that neural network training is either completely low-rank or has multiple phrases with initially high-rank and subsequent low-rank training. For example, Aghajanyan et al. (2021) show that as the model becomes larger or when it is pre-trained for longer, the rank of the update needed to learn a downstream task reduces. Arora et al. (2019) finds that SGD is biased towards low-rank solutions. The existence of Lottery Tickets early in training Frankle et al. (2019) also partially supports this hypothesis, since training a lottery ticket network could effectively be seen as a low-rank approximation to the regular training process. Our empirical analysis (Section 4) shows that pre-trained neural networks exhibit high-rank updates over long trajectories (Figure 4). However, for a sufficiently small trajectory, the training can be effectively approximated by a low-rank update. Given the above results, we speculate that neural network training is locally low-rank, which directly motivates ReLoRA.
@@ 3 Experiments
To evaluate the effectiveness of ReLoRA, we apply it to train a transformer language model on the C4 dataset Raffel et al. (2020) using various model sizes: 60M, 130M, 250M, 350M, and 1.3B.
In all experiments we train without data repetition (single epoch) on at least compute-optimal amount of data, estimated using Chinchilla Scaling Laws Hoffmann et al. (2022).
Algorithm 1 ReLoRA. $\theta$ is model parameters, $\hat{\theta}$ is model parameters with linear layers replaced with ReLoRA, $M$ and $V$ are Adam optimizer states, $\eta$ is learning rate, and $q$ is the reinit frequency.
@@
Params Hidden Heads Layers Learning rate Batch size Seq. len. Data amount
@@
1.3B 2048 24 32 4e-4 786K 512 23.1B
Table 1: Hyperparameters of the language models trained in this study. Batch size and data amount are specified in tokens.
@@
Our architecture is based on transformer Vaswani et al. (2017) and closely resembles LLaMA Touvron et al. (2023). Namely, we use pre-normalization, RMSNorm Zhang and Sennrich (2019), SwiGLU activations Shazeer (2020), $\frac{8}{3}h$ fully-connected hidden state size Touvron et al. (2023), and rotary embeddings Su et al. (2021). We select the number of pre-training tokens based on the Chinchilla scaling laws Hoffmann et al. (2022). Architecture and training hyperparameters are presented in Table 1.
For all LoRA and ReLoRA experiments, we use rank $r=128$ as our initial experiments showed it to have the best perplexity/memory trade-off. You can find additional recommendations on ReLoRA hyperparameter selection in Appendix A. We perform additional experiments comparing different rank choices for the 1.3B model in Section 4.1. We use bfloat16 for all floating point operations and FlashAttention Dao et al. (2022) for effective attention computation.
@@
In our experiments, ReLoRA replaces all attention and fully-connected network parameters, while updating the embeddings and normalization layers full-rank. Since ReLoRA-wrapped models have fewer trainable parameters than full-rank training, we include a Control baseline, which is a full-rank transformer with the same number of trainable parameters as ReLoRA.
We initialize ReLoRA from a checkpoint of full-rank training at 5,000 update steps and reset it every 5,000 steps thereafter, 3 times in total till we reach 20K steps. After each reset, 99% of the optimizer state is pruned based on magnitude, and the loss is warmed up for the next 100 iterations. ReLoRA parameters are reinitialized following LoRA best practices, Kaiming initialization He et al. (2015) for $A$-matrix, and zeros for $B$-matrix.
Scaling up to 1.3B
After initial results at 130M and 350M model sizes, we applied ReLoRA to train a 1.3B parameter language model. As a baseline, we pre-trained a 1.3B model from scratch on 23B tokens. We performed multiple ReLoRA runs starting from 2K, 5K, and 10K checkpoints. In most of the experiments, we continued using $r=128$ and our additional experiments show minimal difference between rank 128 and 512 (hidden size is 2048). Section 4.1 describes these experiments in detail.
@@ 4 Results
Our results are presented in Table 2 and Figure 1. ReLoRA significantly outperforms LoRA training demonstrating the effectiveness of our proposed modifications (ablated in Section 6). Additional pre-training loss figures are available in Appendix C.
Furthermore, ReLoRA achieves similar performance to full-rank training in both upstream and downstream tasks (Table 3). [^2] Note that the absolute values of GLUE results are expected to be quite far from state-of-the-art, because our models were pre-trained on roughly 20 times less data than T5 or BERT.
60M 130M 250M 350M 1.3B
Full training 33.81 (60M) 23.65 (130M) 22.39 (250M) 18.66 (350M) 16.83 (250M)
Control 36.52 (43M) 27.30 (72M) 25.43 (99M) 23.65 (130M) 21.73 (250M)
LoRA 47.44 (43M) 34.17 (72M) 36.60 (99M) 57.11 (125M) -
LoRA + Warm Start 34.73 (43M) 25.46 (72M) 22.86 (99M) 19.73 (125M) 18.23 (250M)
ReLoRA 34.46 (43M) 25.04 (72M) 22.48 (99M) 19.32 (125M) 17.27 (250M)
Training tokens 1.2B 2.6B 6.8B 6.8B 23.1B
Table 2: Language model perplexity when trained using each of the above methods. Number of trainable parameters for each model in (brackets). Control baseline is full-rank training a model with the same total number of parameters as the number of trainable parameters in low-rank training.
CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP Avg
Full-rank pretrained 35.43 83.85 76.96 64.26 88.99 70.98 83.38 84.49 73.54
Not pretrained 7.59 22.73 67.00 51.15 82.61 60.04 67.92 78.40 54.68
ReLoRA 31.07 83.33 78.43 60.65 89.45 72.27 83.93 86.01 73.14
Table 3: Applying ReLoRA to fine-tune 350M models pre-trained full-rank and using ReLoRA. We observe minimal differences between the models.
@@
To determine whether ReLoRA performs a higher rank update than LoRA, we plot the singular value spectrum of the learned update to the warm-start weights. Specifically, the difference between warm-start weights and the final weights for ReLoRA, LoRA, and full-rank trained models. Figure 3 illustrates significant qualitative differences between LoRA and ReLoRA for the singular values of $\Delta W_{Q}$, $\Delta W_{K}$, $\Delta W_{V}$, and $\Delta W_{down}$. While most of the singular values for LoRA are zero (Figure 4) with a noticeable number of exceptionally high values above 1.5, ReLoRA exhibits a higher distribution mass between 0.1 and 1.0, reminiscent of full-rank training.
Additionally, we computed the number of singular values less than 0.1 for LoRA, ReLoRA, and full-rank training. Our results (Figure 4) show that ReLoRA has a much smaller number of near-zero singular values than LoRA, closer to full-rank training. This observation emphasizes the significance of high-rank updates and demonstrates that ReLoRA does accomplish a high-rank update by performing multiple low-rank updates. We also perform ReLoRA component ablation (Table 6) and discuss it in Section 6.
@@
Figure 3: Singular values spectra of the weight difference between ReLoRA and LoRA at 5,000 iterations (warm start) and 20,000 iterations. ReLoRA exhibits a closer resemblance to full-rank training than to LoRA, indicating its effectiveness in approximating full-rank behavior. 350M models.
@@
Figure 4: The number of singular values >0.1 in weight matrices of the learned update. 350M models.
4.1 Scaling up to 1.3B
Figure 5: Training loss at 1.3B scale and the associated baselines. ReLoRA outperforms LoRA throughout training and the gap between the methods increases over time.
1.3B @15K steps 1.3B @20K steps 1.3B @30K steps
Full training 17.67 (250M) 17.00 (250M) 16.83 (250M)
Control 22.67 (250M) 22.00 (250M) 21.73 (250M)
LoRA + Warm Start 18.50 (250M) 18.38 (250M) 18.23 (250M)
ReLoRA 17.94 (250M) 17.64 (250M) 17.27 (250M)
Training tokens (billions) 11.8 15.7 23.1
Table 4: Results at 1.3B scale. Number of trainable parameters for each model in (brackets).
Our best run at this model size starts after a 10K step warm start ($33\%$ of the total update steps). We train ReLoRA with rank $r=128$, learning rate 5e-4, 100 steps lr warmup, and 50 steps restarts warmup. The results are presented in the Figure 5 and Table 4. ReLoRA clearly outperforms LoRA throughout the training with the gap between the methods increasing from 0.56 at 15K steps to 0.96 at 30K steps. At the end of the training, ReLoRA is able to reach a perplexity of 17.24, only 0.44 higher than full-rank training. You can find additional recommendations on ReLoRA hyperparameter selection in Appendix A.
Varying ReLoRA rank
In this experiment we wanted to evaluate if $r=128$ is still applicable to the model of this size (hidden size $2048$) or if it needs to be increased. To do that, we used an early checkpoint for the warm start (5K out of 30K steps). This was beneficial for the comparison, as at this point loss changes quickly which makes any differences in training dynamics more evident. We train these models for additional 10K iterations. Unexpectedly, we found very little difference between ranks 128 (ppl. 19.16) and 512 (ppl. 19.00).
Negative results: Online ReLoRA
250M 1.3B
(@15k steps) (@25k steps)
ReLoRA 27.66 17.36
Online ReLoRA 29.31 17.80
Table 5: Online ReLoRA.
Intuitively, more frequent ReLoRA resets can lead to better performance, as they, in principle, can learn a higher rank update. Usually, for every ReLoRA reset, we would also perform an optimizer reset and learning rate scheduler re-warmup (Section 1). However, in our experiments we observed that very high ReLoRA reset rates lead to worse performance.
Online ReLoRA resolves this issue quite elegantly – it merges LoRA parameters very frequently (e.g., every 100 iterations) while keeping the optimizer reset rate at 2-5K iterations. Unexpectedly, we found that it performs worse than regular ReLoRA at both 250M and 1.3B scales (Table 5).
ReLoRA Training Speedup
Training ReLoRA took 440 A100-hours, saving 56 A100-hours compared to full-rank training. A part of the speedup was due to the ability to use two times larger microbatch size. When training with the same microbatch size, ReLoRA improved RAM consumption from 27.8Gb to 22.3Gb saving 5.5Gb of GPU RAM. Overall, in the 8xA100 setup, combining the warm start and ReLoRA training time, 1.3B-ReLoRA took 86 hours (wall clock) to train compared to 93.5 hours to train 1.3 model full-rank on the same amount of data. This yields a relative speed improvement of $9\%$.
Restarts Optimizer Reset Jagged Schedule Warm Start Perplexity ($\downarrow$)
$\times$ $\times$ $\times$ $\times$ 34.17
$\checkmark$ $\times$ $\times$ $\times$ 34.25
$\checkmark$ $\checkmark$ $\times$ $\times$ (diverged)
$\checkmark$ $\times$ $\checkmark$ $\times$ 34.29
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 29.77
$\times$ $\times$ $\times$ $\checkmark$ 25.46
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 25.04
Regular training 23.65
Table 6: Ablation studies of ReLoRA (130M models). Restarts and warm starts are essential for good performance. Restarts and optimizer resets without a jagged schedule causes the model to diverge.
8xA100 6xA6000 (Ada) 2x3090
Full-rank throughput 137 ex/sec 84 ex/sec 8.8 ex/sec
ReLoRA throughput 157 ex/sec 124 ex/sec 17.8 ex/sec
Immediate speedup 15% 48% 102%
Warm-start adjusted ReLoRA throughput 149 ex/sec 111 ex/sec 14.8 ex/sec
Total speedup 9% 32% 51%
Table 7: Performance metrics in different hardware configurations. Warm start adjustment assumes 33% of full-rank training before switching to ReLoRA.
We additionally observed that ReLoRA speedup is significantly hardware-dependent (Table 7). In our early experiments on 2xRTX3090, we estimated the speedup of $42\%$. In a more practical, but still relatively budget setup of 6xA6000 Ada, we estimated 152 hours of wall-clock training time for the 1B full-rank model and 119 hours for the ReLoRA model with $33\%$ warm start. This saves 33 hours yielding a relative speedup of $21\%$. We attribute the difference to the GPU memory speed. ReLoRA can more effectively utilize low-bandwidth memory as it has less trainable parameters.